OpenStudy (anonymous):

During a given month, P(A)=0.75,P(B)=0.6, and P(A and B)=0.5. Are these two events independent? Are they mutually exclusive? yes;yes yes;no no;yes no;no it cannot be determined

jimthompson5910 (jim_thompson5910):

If event A and event B are independent events, then P(A and B) = P(A) * P(B)

jimthompson5910 (jim_thompson5910):

If event A and event B are mutually exclusive events, then P(A and B) = 0 this can be represented by 2 circles that don't overlap. The lack of the overlapping region means it's impossible to land in both regions at the same time 

OpenStudy (anonymous):

is that no, no?

OpenStudy (anonymous):

o i think no yes is this right

jimthompson5910 (jim_thompson5910):

In this case, is P(A and B) = 0 true?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

ooppps no

jimthompson5910 (jim_thompson5910):

yeah, P(A and B) = 0 is false in this problem since they gave you P(A and B) = 0.5

jimthompson5910 (jim_thompson5910):

how about P(A and B) = P(A) * P(B) ?

OpenStudy (anonymous):

false

jimthompson5910 (jim_thompson5910):

why false?

OpenStudy (anonymous):

um not 100%

jimthompson5910 (jim_thompson5910):

P(A and B) = 0.5 is given what is P(A) * P(B) equal to?

OpenStudy (anonymous):

sorry not sure

OpenStudy (anonymous):

0.6

jimthompson5910 (jim_thompson5910):

P(A)=0.75,P(B)=0.6 so P(A) * P(B) = ??

OpenStudy (anonymous):

0.75

jimthompson5910 (jim_thompson5910):

you just multiply the two, that's what P(A) * P(B) means

OpenStudy (anonymous):

450

jimthompson5910 (jim_thompson5910):

that's way too high

OpenStudy (anonymous):

.45

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so we're given P(A and B) = 0.5 but yet P(A) * P(B) = 0.45 0.5 and 0.45 aren't equal, therefore, P(A and B) = P(A) * P(B) is false so A and B are NOT independent events

jimthompson5910 (jim_thompson5910):

P(A and B) = P(A) * P(B) has to be true for A and B to be independent

OpenStudy (anonymous):

thank you!!

jimthompson5910 (jim_thompson5910):

so somehow, A changes B or B changes A

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

correct answer is actually no;no.......

jimthompson5910 (jim_thompson5910):

yeah that's what I said above